Find the projection of 𝑖 = (4, βˆ’1, 4) on 𝑗 = (βˆ’2, 0, 4).


1
Expert's answer
2021-04-29T14:29:39-0400

Projection of i on j "=\dfrac{\vec{i}.\vec{j}}{|\vec{j}|}"


"=\dfrac{(4\hat{i}-\hat{j}+4\hat{k}).(-2\hat{i}+4\hat{k})}{\sqrt{(-2)^2+(4)^2}}"


"=\dfrac{-8+16}{\sqrt{4+16}}=\dfrac{8}{\sqrt{20}}=\dfrac{4\sqrt{5}}{5}"